# Composable Finite-Size Security of High-Dimensional Quantum Key Distribution Protocols

### v1

@@ III.2 Lift to coherent attacks
It remains to lift the proof to coherent attacks, based on the postselection technique by Refs. [37, 38]. We follow the notation by [38]. This requires some additional notation. Recall that $\Phi:~{}ABE\mapsto K_{A}K_{B}E$, where Eve’s register $E$ contained all information she gathered during the protocol execution and define $\Phi^{(\ell)}:=\mathrm{Tr}_{E}\circ\mathrm{Tr}_{K_{B}}\circ\Phi$, a QKD protocol map that produces a key of length $\ell$ upon acceptance. Additionally, let $\Phi^{(\ell)}_{\mathrm{ideal}}$ be the corresponding ideal QKD protocol map. Now, in contrast to the first part of the proof, we now assume a general $\rho_{ABE}=\rho_{A^{n}B^{n}E^{n}}$ without any known structure, so in particular $\rho_{ABE}\neq\tau_{ABE}^{\otimes n}$. After splitting off the correctness part, similar to the argument above, the secrecy statement reads
$\frac{1}{2}\left|\left|\left(\left(\Phi^{(\ell)}-\Phi_{\mathrm{ideal}}^{(\ell)}\right)\otimes\mathrm{id}_{E}^{n}\right)\left(\rho_{A^{n}B^{n}E^{n}}\right)\right|\right|_{1}\leq\epsilon_{\mathrm{sec}}.$ (6)
Since we applied a random permutation to the classical data acquired during the protocol execution, for measurements that act only on single rounds, the considered protocol is permutation invariant such that the map $\Phi^{(\ell)}-\Phi_{\mathrm{ideal}}^{(\ell)}$ is permutation invariant, too. For such maps, the postselection theorem [37, Theorem 1] gives an upper bound on this expression in terms of a purification of so-called de Finetti states.
This allows us to relate the $\epsilon$-secrecy of the QKD protocol for an arbitrary state $\rho_{ABE}$ to the $g_{n,x}\epsilon$-secrecy of the same protocol if the input is a purification of a probabilistic mixture of i.i.d. states. The final step of the postselection method then comprises of relating this to the secrecy of an i.i.d. state. The (corrected) postselection Theorem (Theorem 3 in [38]) states the following; Providing that one has proven security of a QKD protocol for i.i.d. states, yielding a key of length $\ell$
$\frac{1}{2}\left|\left|\left(\left(\Phi^{(\ell)}-\Phi_{\mathrm{ideal}}^{(\ell)}\right)\otimes\mathrm{id}_{E}^{n}\right)\left(\tau_{ABE}^{\otimes n}\right)\right|\right|_{1}\leq\epsilon_{\mathrm{sec}}$
with security parameter $\epsilon_{\mathrm{sec}}=\max\{\epsilon_{\mathrm{AT}},\epsilon_{\mathrm{PA}}\}$ (which is what we have just accomplished before), the security of the same protocol with a purification of a de Finetti state as input that hashes to a length $\ell^{\prime}=\ell-2\log_{2}\left(g_{n,x}\right)$ then is
$\frac{1}{2}\left|\left|\left(\left(\Phi^{(\ell^{\prime})}-\Phi_{\mathrm{ideal}}^{(\ell^{\prime})}\right)\otimes\mathrm{id}_{R}\right)\left(\sigma_{A^{n}B^{n}R}\right)\right|\right|_{1}\leq\epsilon_{\mathrm{sec}}^{\prime}$ (7)
secure, with security parameter $\epsilon_{\mathrm{sec}}^{\prime}=\epsilon_{\mathrm{PA}}+2\sqrt{2\epsilon_{\mathrm{AT}}}$. Note that while the new postselection theorem by Ref. [38] straightforwardly extends from prepare-and-measure protocols to entanglement-based protocols
@@
Let $\epsilon_{\mathrm{EV}},\epsilon_{\mathrm{PA}},\epsilon_{\mathrm{AT}},\bar{\epsilon}>0$. Let $N$ be the total number of quantum signals exchanged, $k_{T}$ the number of test rounds, $n:=N-k_{T}$. The objective general high-dimensional QKD protocol is $\epsilon_{\mathrm{EV}}+g_{n,x}\left(\epsilon_{\mathrm{PA}}+2\sqrt{2\epsilon_{\mathrm{AT}}}\right)$-secure against coherent attacks, if, in case the protocol does not abort, the key is hashed to a length of
@@
In this work, we use $g_{n,x}\leq\left(\frac{e(n+x-1)}{x-1}\right)^{x-1}$ from Ref. [38] as an upper bound. Summing up, based on the security of our i.i.d. collective attacks proof, we now can establish security against coherent attacks for a shorter key with a modified security parameter. When we present our results, we plot secure key rates instead of achievable secure key lengths, thus we divide both sides by $N$ and obtain
@@ IV Variable-Length Security
In the previous section, we have proven security of a high-dimensional QKD protocol that produces key of fixed length, or aborts. The most crucial part of such a fixed-length protocol in reality is to construct an acceptance set, which practically boils down to defining the expected channel behavior ($r_{X}$, potentially along with some margin $t_{X}$). This is because in many practical applications, the expected channel behaviour is not known or can change rapidly, such as in free-space and satellite-based QKD applications, which are two of the prototypical use-cases where high-dimensional QKD can play out its strengths. However, according to the security argument given in Section III, the accepted statistics have to be fixed before the run of the protocol and the abort decision is made by comparing the actual measurement results with the acceptance set. For unknown or rapidly changing values, this would mean that even protocols with non-unique acceptance abort very often (and unique acceptance protocols basically all the time), leading to a low expected key rate [41, 39]. The only possible counter-action is increasing the parameters $t_{X}$ to enlarge the acceptance set. This, however, diminishes the secure key length, as the reported length has to be compatible with the worst-case state within the given set, while not increasing the acceptance probability for varying channels significantly.
We address this problem by introducing a variable-length version of our protocol and proving security, applying the theory by Ref. [39]. We start by stating the required protocol modifications and introducing some additional notation.
@@
For the fixed-length version of the protocol the value of $\ell_{i}$ and $\lambda_{i}^{\mathrm{EC}}$ had to be fixed before the execution of the protocol, now this decision is made based upon the data. Let $\vec{F}^{\mathrm{obs}}\in\mathbb{R}^{|\Theta|}$ denote the vector containing the experimental observations for all considered observables $X\in\Theta$ and denote $\Phi_{X}(\rho):=\mathrm{Tr}\left[\rho X\right]$.
@@ IV.1 Security against collective attacks
In what follows, we summarize the idea of the variable-length security proof from Ref.  [39] and adapt it to the objective high-dimensional QKD protocol. First, based on their observations $\vec{F}^{\mathrm{obs}}$, Alice and Bob need to find a statistical estimator $b_{\mathrm{stat}}\left(\vec{F}^{\mathrm{obs}}\right)$, which is a high-probability lower bound to the Rényi entropy $H_{\alpha}\left(Z^{n}|C^{n}E^{n}\right)_{\rho}$ of the underlying quantum state with high probability, where $C$ denotes the register collecting the leaked classical information,
@@
This requires the construction of a set $V\left(\vec{F}^{\mathrm{obs}}\right)$ such that $\mathrm{Pr}\left[\tau_{AB}\in V\left(\vec{F}^{\mathrm{obs}}\right)\right]\geq 1-\epsilon_{\mathrm{AT}}$. Based on this estimator and the public announcements, Alice and Bob determine the information leakage $\lambda^{\mathrm{EC}}\left(\vec{F}^{\mathrm{obs}}\right)$, which is some integer number between 00 and $\lambda_{\mathrm{max}}^{\mathrm{EC}}$. This allows calculating the secure key length as
$\ell\left(\vec{F}^{\mathrm{obs}}\right):=$ (11)
@@
In the fixed-length scenario, the acceptance test was built upon defining expected statistics and finding the set of density matrices that could have produced these statistics only with some small probability; the complement of this set was the set $\mathcal{S}_{\mathrm{AT}}$. In the variable-length scenario, the construction of the acceptance set follows a slightly different logic. A fixed but otherwise unknown quantum state gives rise to observed statistics $\vec{F}^{\mathrm{obs}}$, which is a random variable. Based on this observation, we aim to construct a set $V(\vec{F}^{\mathrm{obs}})$ such that $\mathrm{Pr}\left[\tau_{AB}\in V(\vec{F}^{\mathrm{obs}})\right]\geq 1-\epsilon_{\mathrm{AT}}$. We are using observables, thus we have to slightly adapt the argument from [39]. Let $\Theta$ denote the set of observables used in the protocol and let $x:=||X||_{\infty}=||X||_{\infty}||\rho||_{1}\geq||X\rho||_{1}$ for $X\in\Theta$. For each $X\in\Theta$, Hoeffding’s inequality [46] tells us
@@
$V\left(\vec{F}^{\mathrm{obs}}\right):=\left\{\sigma\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{E}):\right.$ (16)
@@
Note that in the present case despite the different argument, the main difference to the acceptance set for the fixed-length scenario is the decomposition of $\epsilon_{\mathrm{AT}}$ as sum of the single-observable epsilons $\epsilon_{X}$. This, however, is due to using Hoeffding’s inequality and can differ more significantly for different sampling arguments.
@@
$b_{\mathrm{stat}}\left(\vec{F}^{\mathrm{obs}}\right):=$ $n\min_{\tau_{AB}\in\mathcal{S}^{\mathrm{AT}}\left(\vec{F}^{\mathrm{obs}}\right)}H_{\mathrm{min}}\left(X|E\right)_{\Phi_{\mathrm{var}}\left(\tau_{ABE}\right)}$ (17)
@@
$\ell\left(\vec{F}^{\mathrm{obs}}\right):=$ (18)
@@ IV.2 Security of variable-length HD-QKD protocols against coherent attacks
$\frac{1}{2}\left|\left|\left(\Phi_{\mathrm{var}}^{(\ell_{1},...,\ell_{M})}-\Phi_{\mathrm{var,ideal}}^{(\ell_{1},...,\ell_{M})}\right)\left(\rho_{A^{n}B^{n}E^{n}}\right)\right|\right|_{1}\leq\epsilon_{\mathrm{PA}}+\epsilon_{\mathrm{AT}},$ (19)
and $\epsilon_{\mathrm{EV}}$-correct. By a similar argument as in the previous section, exploiting the permutation invariance of the objective high-dimensional QKD protocol (see Ref. [38] for details), one can show that providing the protocol is secure against collective attacks with security parameter $\epsilon_{\mathrm{PA}}+\epsilon_{\mathrm{AT}}$, the protocol that hashes keys to length $\ell_{i}^{\prime}:=\ell_{i}-2\log_{2}(g_{n,x})-2\log_{2}\left(\frac{1}{\tilde{\epsilon}}\right)$ instead of length $\ell_{i}$ upon $\Omega_{i}$ with security parameter $g_{n,x}\left(\sqrt{8\left(\epsilon_{\mathrm{PA}}+\epsilon_{\mathrm{AT}}\right)}+\frac{\tilde{\epsilon}}{2}\right)$,
$\frac{1}{2}\left|\left|\left(\Phi_{\mathrm{var}}^{(\ell_{1}^{\prime},...,\ell_{M}^{\prime})}-\Phi_{\mathrm{var,ideal}}^{(\ell_{1}^{\prime},...,\ell_{M}^{\prime})}\right)\left(\rho_{A^{n}B^{n}E^{n}}\right)\right|\right|_{1}$ (20)
@@
$\ell\left(\vec{F}^{\mathrm{obs}}\right):=$ (22)
$-2\log_{2}\left(g_{n,x}\right)-2\log_{2}\left(\frac{1}{\tilde{\epsilon}}\right),$
@@ V Security proof method
Having discussed the security for two different attack scenarios and both the fixed and variable-length scenarios, it remains to find a way to evaluate the given expressions. However, as discussed earlier and found in Ref. [27], widely available convex optimization methods [30, 29] cannot handle high-dimensional QKD protocols. On the other hand, computationally less demanding methods [22, 24] rely on hard-to-implement measurements, which for most practical setups are not available. Instead, we employ a recent technique developed in Ref. [31] that provides lower bounds on the secure key rate without high computational demands, solely relying on available measurements.
@@ V.1 Key rate evaluation
While this method computes asymptotic key rates, the formulation of our composable finite-size key rate finding problems derived in Sections III and IV in terms of the min-entropy already allows us to directly apply this framework. The basic idea of this method is the following [41]. It is well known that the min-entropy can be expressed as the negative binary logarithm of Eve’s probability of guessing Alice’s key string correctly, $H_{\mathrm{min}}(X|E)_{\rho}=-\log_{2}\left(p_{\mathrm{guess}}\right)$. This allows us to recast the minimization of the min-entropy over $\mathcal{S}^{\mathrm{AT}}$ as maximization over all possible measurements Eve might carry out in order to guess Alice’s outcomes. Applied to our finite-size optimization problem, this turns into
@@
$~{}~{}~{}~{}~{}~{}~{}y_{0}\geq\lambda_{\mathrm{max}}\left(M_{\ell}\right)~{}~{}~{}\forall\ell=0,...,d-1$
$~{}~{}~{}~{}~{}~{}~{}z_{0}^{L},z_{0}^{U},z_{j}\geq 0~{}~{}~{}\forall j=1,...,N_{W}$
@@
which boils down the problem to solve a rather simple minimization problem subject to a constraint that depends on the maximal eigenvalue of a matrix which is a function of the chosen observables $M_{\ell}:=\outerproduct{\ell}{\ell}\otimes\mathbbm{1}_{d}-\left(z_{0}^{U}-z_{0}^{L}\right)\bar{W}_{0}-\sum_{j=1}^{N_{W}}z_{j}^{U}\bar{W}_{j}$.
Thus, the crucial task was reduced to finding experimentally accessible observables. This can be significantly simplified by extending the set of accessible operators using a matrix-completion technique [47, 20]. For further details, we refer the reader to Ref. [31].

### v2

@@ III.2 Lift to coherent attacks
It remains to lift the proof to coherent attacks, based on the postselection technique by Refs. [Christandl_2009, Nahar_2024]. We follow the notation used in Ref. [Nahar_2024]. This requires some additional notation. Recall that $\Phi:~ABE\mapsto K_{A}K_{B}E$, where Eve’s register $E$ contains all information she gathered during the protocol execution and define $\Phi^{(\ell)}:=\mathrm{Tr}_{E}\circ\mathrm{Tr}_{K_{B}}\circ\Phi$, a QKD protocol map that produces a key of length $\ell$ upon acceptance. Additionally, let $\Phi^{(\ell)}_{\mathrm{ideal}}$ be the corresponding ideal QKD protocol map. Now, in contrast to the first part of the proof, we now assume a general $\rho_{ABE}=\rho_{A^{n}B^{n}E^{n}}$ without any known structure, so in particular $\rho_{ABE}\neq\tau_{ABE}^{\otimes n}$. After splitting off the correctness part, similar to the argument above, the secrecy statement reads
[…]
Since we applied a random permutation to the classical data acquired during the protocol execution, for measurements that act only on single rounds, the considered protocol is permutation invariant such that the map $\Phi^{(\ell)}-\Phi_{\mathrm{ideal}}^{(\ell)}$ is also permutation invariant (for a formal definition of permutation invariance of a map, we refer to [Christandl_2009, Nahar_2024]). For such maps, the postselection theorem [Christandl_2009, Theorem 1] gives an upper bound on this expression in terms of a purification of so-called de Finetti states.
This allows us to relate the $\epsilon$-secrecy of the QKD protocol for an arbitrary state $\rho_{ABE}$ to the $g_{n,x}\epsilon$-secrecy of the same protocol if the input is a purification of a probabilistic mixture of i.i.d. states. The final step of the postselection method then comprises relating this to the secrecy of an i.i.d. state. The (corrected) postselection Theorem (Corollary 3.1 in [Nahar_2024]) states the following: Providing that one has proven security of a QKD protocol for i.i.d. states, yielding a key of length $\ell$
[…]
with security parameter $\epsilon_{\mathrm{sec}}=\max\{\epsilon_{\mathrm{AT}},\epsilon_{\mathrm{PA}}\}$ (which is what we have just accomplished before), the security of the same protocol with a purification of a de Finetti state as input that hashes to a length $\ell^{\prime}=\ell-2\log_{2}\left(g_{n,x}\right)$, where $g_{n,x}={n+x-1\choose n}$, is then
[…]
secure, with security parameter $\epsilon_{\mathrm{sec}}^{\prime}=\epsilon_{\mathrm{PA}}+2\sqrt{2\epsilon_{\mathrm{AT}}}$. Note that the new postselection theorem by Ref. [Nahar_2024] straightforwardly extends from prepare-and-measure protocols to entanglement-based protocols.
@@
Let $\epsilon_{\mathrm{EV}},\epsilon_{\mathrm{PA}},\epsilon_{\mathrm{AT}},\bar{\epsilon}>0$. Let $N$ be the total number of quantum signals exchanged, $k_{T}$ be the number of test rounds, and $n:=N-k_{T}$. The objective general high-dimensional QKD protocol is $\epsilon_{\mathrm{EV}}+g_{n,x}\left(\epsilon_{\mathrm{PA}}+2\sqrt{2\epsilon_{\mathrm{AT}}}\right)$-secure against coherent attacks, if, in case the protocol does not abort, the key is hashed to a length of
@@
In this work, we use $g_{n,x}\leq\left(\frac{e(n+x-1)}{x-1}\right)^{x-1}$ from Ref. [Nahar_2024] as an upper bound. In non-technical terms, Theorem 3 allows us to relate the security against general attacks to the security against i.i.d. collective attacks, which we established earlier in Theorem 2. While we still can relate the achievable secure key length $\ell^{\prime}$ to a similar optimisation problem as we found for collective attacks, the stronger security claim comes at a price. In order to obtain secure key, the additional correction term $\log_{2}\left(g_{n,x}\right)$ tells us that we need to hash our raw key a bit more, while additionally, the total security parameter grows.
When we present our results, we plot secure key rates instead of achievable secure key lengths; thus, we divide both sides by $N$ and obtain
@@ IV Variable-Length Security
In the previous section, we have proven the security of a high-dimensional QKD protocol that produces a key of fixed length or aborts. The most crucial part of such a fixed-length protocol in reality is to construct an acceptance set, which practically boils down to defining the expected channel behavior ($r_{X}$, potentially along with some margin $t_{X}$). This is because in many practical applications, the expected channel behaviour is not known or can change rapidly, such as in free-space and satellite-based QKD applications, which are two of the prototypical use-cases where high-dimensional QKD can play out its strengths. However, according to the security argument given in Section III, the accepted statistics have to be fixed before the run of the protocol, and the abort decision is made by comparing the actual measurement results with the acceptance set. For unknown or rapidly changing values, this would mean that even protocols with non-unique acceptance abort very often (and unique acceptance protocols basically all the time), leading to a low expected key rate [Kanitschar_2023, Tupkary_2024]. The only possible counter-action is increasing the parameters $t_{X}$ to enlarge the acceptance set. This, however, diminishes the secure key length, as the reported length has to be compatible with the worst-case state within the given set, while not increasing the acceptance probability for varying channels significantly.
We address this problem by introducing a variable-length version of our protocol and proving security, applying the theory by Ref. [Tupkary_2024]. We start by stating the required protocol modifications and introducing some additional notation.
@@
For the fixed-length version of the protocol, the value of $\ell_{i}$ and $\lambda_{i}^{\mathrm{EC}}$ had to be fixed before the execution of the protocol; now this decision is made based upon the data. Let $\vec{F}^{\mathrm{obs}}\in\mathbb{R}^{|\Theta|}$ denote the vector containing the experimental observations for all considered observables $X\in\Theta$ and denote $\Phi_{X}(\rho):=\mathrm{Tr}\left[\rho X\right]$.
@@ IV.1 Security against collective attacks
In what follows, we summarize the idea of the variable-length security proof from Ref.  [Tupkary_2024] and adapt it to the objective high-dimensional QKD protocol. First, based on their observations $\vec{F}^{\mathrm{obs}}$, Alice and Bob need to find a statistical estimator $b_{\mathrm{stat}}\left(\vec{F}^{\mathrm{obs}}\right)$, which is a high-probability lower bound to the Rényi entropy $H_{\alpha}\left(Z^{n}|C^{n}E^{n}\right)_{\rho}$ of the underlying quantum state with, where $C$ denotes the register collecting the leaked classical information,
@@
Intuitively, this estimator provides a high probability lower-bound on Eve’s information about the key, based on the observed statistics $\vec{F}^{\mathrm{obs}}$. Finding such an estimator requires the construction of a set $V\left(\vec{F}^{\mathrm{obs}}\right)$ that contains the (unknown) quantum state responsible for Alice’s and Bob’s observations with high probability, i.e., $\mathrm{Pr}\left[\tau_{AB}\in V\left(\vec{F}^{\mathrm{obs}}\right)\right]\geq 1-\epsilon_{\mathrm{AT}}$. Based on this estimator and the public announcements, Alice and Bob determine the information leakage $\lambda^{\mathrm{EC}}\left(\vec{F}^{\mathrm{obs}}\right)$, which is an integer number between 0 and $\lambda_{\mathrm{max}}^{\mathrm{EC}}$ quantifying how many bits of the key Eve learns by listening to Alice’s and Bob’s communication over the classical channel. An illustration of the variable-length decision can be found in Figure 1.
This allows calculating the secure key length as the difference between the statistical estimator and the error correction leakage, and some additional correction terms of lower order
$\ell\left(\vec{F}^{\mathrm{obs}}\right)$= (11)
@@
[…]
@@
$V\left(\vec{F}^{\mathrm{obs}}\right)=\left\{\sigma\in\mathcal{D}(\mathcal{H}_{A}\otimes\mathcal{H}_{B}\otimes\mathcal{H}_{E}):\right$. (16)
@@
Note that in the present case, despite the different argument, the main difference to the acceptance set for the fixed-length scenario is the decomposition of $\epsilon_{\mathrm{AT}}$ as a sum of the single-observable epsilons $\epsilon_{X}$. This, however, arises from using Hoeffding’s inequality and can differ more significantly for different sampling arguments.
@@
$b_{\mathrm{stat}}\left(\vec{F}^{\mathrm{obs}}\right)$= $n\min_{\tau_{AB}\in V\left(\vec{F}^{\mathrm{obs}}\right)}H_{\mathrm{min}}\left(X|E\right)_{\Phi_{\mathrm{var}}\left(\tau_{ABE}\right)}$ (17)
@@
$\ell\left(\vec{F}^{\mathrm{obs}}\right)$= (18)
@@
Intuitively, Theorem 4 states that we can obtain a qualitatively similar security claim as we obtained in the fixed-length scenario (see Theorem 2) even if we construct our statistical analysis around the observed rather than the expected measurement outcomes. However, as outlined before the statement, this requires a different security argument and leads to different correction terms.
@@ IV.2 Security of variable-length HD-QKD protocols against coherent attacks
[…]
@@
$\ell\left(\vec{F}^{\mathrm{obs}}\right)$= (22)
$\max\left\{0,b_{\mathrm{stat}}\left(\vec{F}^{\mathrm{obs}}\right)-\lambda^{\mathrm{EC}}\left(\vec{F}^{\mathrm{obs}}\right)-\theta(\alpha,\epsilon_{\mathrm{PA}},\epsilon_{\mathrm{EV}})\right$.
[…]
@@
Similar to the variable-length statement for collective attacks, Theorem 5 presents a variable-length version for the achievable secure key length against coherent attacks, which builds the statistical analysis around the observed rather than the expected behaviour of the channel. The core message is that we obtain a qualitatively similar security statement to the fixed-length case, even if we build our security argument upon the observed rather than the expected statistics.
Figure 1: Flowchart illustrating the variable-length decision. The length of the key is now a function of the observed statistics $\vec{F}^{\mathrm{obs}}$.
@@ V Security proof method
Having discussed the security for two different attack scenarios and both the fixed and variable-length scenarios, it remains to find a way to evaluate the given expressions. However, as discussed earlier and found in Ref. [BKPH_2023], widely available convex optimization methods [Araujo_2022, Winick_2018] cannot handle high-dimensional QKD protocols. On the other hand, computationally less demanding methods [Sheridan_2010, Doda_2021] rely on hard-to-implement measurements, which for most practical setups are not available. Instead, we employ a recent technique developed in Ref. [Kanitschar_2024a] that provides lower bounds on the secure key rate without high computational demands, solely relying on available measurements.
@@ V.1 Key rate evaluation
While this method computes asymptotic key rates, the formulation of our composable finite-size key rate finding problems derived in Sections III and IV in terms of the min-entropy already allows us to directly apply this framework. The basic idea of this method is the following [Kanitschar_2023]. It is well known that the min-entropy can be expressed as the negative binary logarithm of Eve’s probability of guessing Alice’s key string correctly, $H_{\mathrm{min}}(X|E)_{\rho}=-\log_{2}\left(p_{\mathrm{guess}}\right)$. This allows us to recast the minimization of the min-entropy over $\mathcal{S}^{\mathrm{AT}}$ as maximization over all possible measurements Eve might carry out in order to guess Alice’s outcomes. Applied to our finite-size optimization problem, this turns into
@@
$~~~~~~~y_{0}\geq\lambda_{\mathrm{max}}\left(M_{\ell}\right)~~~\forall\ell=0,.,d-1$
$~~~~~~~z_{0}^{L},z_{0}^{U},z_{j}\geq 0~~~\forall j=1,.,N_{W}$
@@
[…]
Thus, the crucial task was reduced to finding experimentally accessible observables. This can be significantly simplified by extending the set of accessible operators using a matrix-completion technique [Tiranov_2017, Martin_2017]. For further details, and a discussion about the complexity scaling, we refer the reader to Ref. [Kanitschar_2024a].